On January 1, 2022, Sage Hill Inc. had these stockholders’ equity balances.

Common Stock, $1 par (2,500,000 shares authorized, 650,000 shares issued and outstanding) $650,000
Paid-in Capital in Excess of Par Value 1,420,000
Retained Earnings 698,000
Accumulated Other Comprehensive Income 50,000


During 2022, the following transactions and events occurred.

1. Issued 46,500 shares of $1 par value common stock for $2 per share.
2. Issued 61,000 shares of common stock for cash at $6 per share.
3. Purchased 22,500 shares of common stock for the treasury at $3.05 per share.
4. Declared and paid a cash dividend of $205,500.
5. Earned net income of $422,000.
6. Had other comprehensive income of $15,800.


Prepare the stockholders’ equity section of the balance sheet at December 31, 2022.
Answer #1
SAGE HILL INC
Partial balance sheet
stockholders' equity
paid in capital
common stock 757500
additional paid in capital
paid in capital in excess of par value -common stock 279000
total paid in capital 1036500
Retained earnings 914500
total paid in capital & retained earning 1951000
Other comprehensive income 65800
less:treasury stock (22500*3.05) -68625
total stockholder equity 1948175
Working
common stock (650000+46500+61000)*1 = 757500
Paid in capital in excess of par (1420000+46500*1+61000*5)=279000
Retained earnings (698000+422000-205500) = 914500
other coprehensive income (50000+15800) = 65800

Please give a thumbs up if it is helpful & let me know if any doubt